Give the coordinates of the x-intercept(s)(if they exist) for the graph of the equation: y=^2+6x+9
Answer by waynest(281)

y = x^2 + 6x + 9
y = x^2 + 6x + (6/2)^2 - (6/2)^2 + 9
y = X62 + 6x + 9 - 9 + 9
y = (x + 3)^2
x-intercept = (-3, 0)